A university employs a total of 420 teaching assistants and research assistants. There are two times as many teaching assistants as research assistants. Find the number of research assistants employed by the university.
research assistants
1
Expert's answer
2020-10-05T17:09:13-0400

Answer :

let

r= research assistants

t= teaching assistants

given , r+t=420

and t= 2r

so substitute t=2r in the equation r+t=420

r+2r=420

3r=420

r=420/3

r=140

t=2*r

t= 2*140

t =280

so the number of research assistants employed by the university are 140